Christine's Chronographs makes two models of a high-end watch—the model 94 and the model 101. Both models go through a special polishing step that requires the use of the single polishing machine at the company. The price and costs of the two models appear as follows.

Model 94 Model 101
Price per watch $ 5,100 $ 9,700
Variable cost per watch $ 2,600 $ 5,800
Polishing hours per watch 2 3

The polishing machine used for both models has a capacity of 33,600 hours annually. Fixed manufacturing costs are $7,800,000 annually.

Required:

a. Suppose that the maximum annual unit sales that Christine's can achieve is 14,000 units of the Model 94 and 12,200 of the Model 101. How many units of each watch model should Christine's produce annually?

b. Suppose that the maximum annual unit sales that Christine's achieve make is 12,000 units of the Model 94 and 9,000 of the Model 101. How many units of each watch model should Christine's produce annually?

2..

Freeflight Airlines is presently operating at 70 percent of capacity. Management of the airline is considering dropping Freeflight's routes between Europe and the United States. If these routes are dropped, the revenue associated with the routes would be lost and the related variable costs saved. In addition, the company's total fixed costs would be reduced by 20 percent.

Segmented income statements for a typical month appear as follows (all amounts in millions of dollars):

Routes Within U.S. Within Europe Between U.S.
and Europe
Sales $ 3.28 $ 2.77 $ 2.82
Variable costs 1.26 0.91 1.71
Fixed costs allocated to routes 1.68 1.27 1.37
Operating profit (loss) $ 0.34 $ 0.59 $ (0.26 )

Required:

a. Prepare a differential cost schedule.
Answer #1
Model 94 Model 101
Sales Price $             5,100 $              9,700
Variable Cost $             2,600 $              5,800
Contribution Margin per unit $             2,500 $              3,900
Polishing hours per unit 2 3
Contribution Margin per hour $             1,250 $              1,300
Maximum Demand 14000 12200
Hours required for demand 28000 36600

a. Since contribution margin per hour is higher for Model 101, so polishing hours will first be used for Model 101, and any remaining hours if any for Model 94

Model 101 = 33600 / 3 = 11200 units
Model 94 = 0 units

b.

Model 94 Model 101
Maximum Demand 12000 9000
Hours required for demand 24000 27000


Model 101 = 9000 units
Model 94 = (33600-27000) / 2 = 3300 units

As per HOMEWORKLIB RULES, we can answer only 1 question, i have answered 1, so kindly post others separately.

If you have any query, kindly comment with your query and please mark thumbs up.